OpenStudy (anonymous):

the line parallel to y=-4/3x+1 with a y intercept of 3 please help i dont understand

OpenStudy (ranga):

Parallel lines have the same slope. What is the slope of the line y=-4/3x+1 ?

OpenStudy (anonymous):

the slope would be -4/3

OpenStudy (ranga):

Yes. And so the parallel line will also have the slope of -4/3 The general equation of a line in slope-intercept form is: y = mx + b where m is the slope and b is the intercept. For the parallel line, you already have slope m = -4/3 They tell you the parallel line has the intercept 3 which means b = 3 Just plug them into y = mx + b and you will have the equation of the parallel line.

OpenStudy (anonymous):

y=-4/3x+3?

OpenStudy (ranga):

You got it!

OpenStudy (anonymous):

thanks that was pretty easy lol

OpenStudy (ranga):

you are welcome. It is easy once you get familiar with it.